# Discovering Interpretable Programmatic Policies via Multimodal LLM-assisted Evolutionary Search

### v1

@@ 3.4 Evolutionary Search for Discovering Policy
This subsection details the evolutionary search loop in MLES. Building upon the EoH framework (Liu et al. 2024a), we integrate behavioral evidence analysis, thereby introducing a prototype implementation of MLES designed to discover high-performing programmatic policies. At each search step, parent selection determines the starting points, while evolutionary operators guide the sampling of new policies. By combining carefully designed operators with principled selection strategies, MLES achieves a dynamic balance between exploration and exploitation, facilitating the search for high-performing policies within the policy space.
Evolutionary operators.
The core of MLES is a set of evolutionary operators that drive the generation of new policies. We define four operators, categorized into two types:
@@
Multimodal modification operators (M1_M and M2_M): These operators instruct MLLMs to analyze IBE and refine existing policies through detailed, informed modification. M1_M asks the MLLMs to identify behavioral shortcomings by jointly analyzing the policy code and its IBE, then revise the control logic accordingly. M2_M prompts the MLLMs to identify critical parameters in the policy and adjust them based on the observed evidence, leading to targeted optimization. The integration of IBE analysis provides grounded insights for the MLLM’s reasoning, resulting in more coherent and purposeful offspring generation.
Parent selection.
For each operator, a subset of candidate policies is selected from the policy pool to serve as parents. The selection follows a rank-based probabilistic scheme common in evolutionary computation. Specifically, the selection probability of the $i$-th policy is defined as $p_{i}\propto 1/(r_{i}+N)$, where $r_{i}$ denotes the rank of the policy (according to quantitative metrics), and $N$ is the size of the policy pool. This scheme encourages the exploration of elite policies while maintaining diversity by occasionally selecting lower-ranked individuals.
Few-shot prompt construction.
After selecting the operator and parents, the prompt sampler constructs a task-specific few-shot prompt to guide the MLLMs in generating new policies. Each prompt consists of the following components: (1) a task description, (2) the code and thought of each parent policy, (3) relevant IBE, and (4) operator-specific instructions. Together, these elements provide the necessary context and guidance for effective policy generation.
Policy generation & individual construction.
Guided by the constructed prompts, the LLM ensemble generates multiple candidate policies in parallel. These policies are then executed by evaluators to assess their quantitative metrics. The raw data from these executions is processed by the behavior summarizer, converting it into IBE, and ultimately leading to the construction of the corresponding individual.
Policy pool management.
Offspring individuals are filtered for redundancy and ranked based on their quantitative metrics. The top-performing and diverse candidates are then added back into the policy pool, enabling progressive refinement and sustained diversity over successive iterations.
@@ 4 Experiments
We conduct empirical experiments to evaluate the effectiveness, efficiency, interpretability, and generalization capabilities of MLES.
@@ 4.1 Experimental Setting
Benchmarks.
We evaluate the proposed MLES framework on two representative control tasks from the OpenAI Gym suite (Brockman et al. 2016): Lunar Lander and Car Racing. These tasks collectively cover both discrete and continuous control settings, providing a comprehensive testbed for our method.
  • Lunar Lander is a discrete-action control task where the agent must land a spacecraft on a designated pad. The agent has four possible discrete actions: doing nothing, applying thrust from the main engine, or firing the left/right orientation engines. The environment has an 8-dimensional state space and is commonly used for benchmarking RL methods in discrete domains.
@@
Baselines.
We compare MLES against both DRL methods and an existing LES approach for direct policy discovery:
  • Deep Q-Network (DQN) (Mnih et al. 2013) is a value-based DRL algorithm designed for discrete action spaces. It serves as a representative DRL baseline for Lunar Lander.
  • Proximal Policy Optimization (PPO) (Schulman et al. 2017) is a widely used on-policy policy gradient method that can handle both discrete and continuous control tasks.
  • EoH (Liu et al. 2024a) is a recent LES framework for automated code discovery. We use it as a baseline to assess the added value of behavioral evidence analysis in MLES, as it shares the same evolutionary framework but lacks the visual feedback-driven behavior analysis used in MLES.
Implementation Details.
For the Lunar Lander task, we select five representative instances for training, while for the Car Racing task, four representative tracks are chosen. Both tasks use ten test environments, each generated from a random seed in the range zero to nine. The quantitative metric used for evaluating performance in the Lunar Lander task is the normalized weight score (NWS), defined as:
where $R$ is the mean episode reward, $C$ is the mean fuel consumption, and $S$ is the success rate of completing the landing task, all computed across multiple instances. For the Car Racing task, we use the average track completion rate across instances as the quantitative metric. For DRL methods, we adopt the default reward function provided by Gym. Further details on these metrics are available in Appendix B.
Both MLES and EoH employ GPT-4o-mini for policy generation. The LLM query budget is set to 2000 requests, corresponding to 10,000 and 8,000 environment resets for Lunar Lander and Car Racing, respectively. To ensure a fair comparison, we train all baselines under an identical budget of environment resets. The evolutionary search population size is set to 16, with the initial population of policies for MLES shared with EoH to avoid any biases arising from different starting points. All experiments are conducted on a machine equipped with an Intel Core i9-13980HX processor and 32 GB of RAM. Each experiment is repeated five times to ensure the robustness and reliability of the results.
@@ 4.2 Experimental Results
Train Test Train Test
DQN 1.017 0.508 79.777 71.724
PPO 1.032 0.846 99.212 94.546
@@
EoH 1.053 0.776 89.808 79.286
MLES 1.090 0.819 98.704 96.358
Table 1: Performance on benchmarks (5 runs). Best results are in bold, second-best are underlined. Higher is better.
Comparative evaluations of MLES
@@
Table 1 reports the average performance across five independent runs for each method on both training and testing instances. MLES gets the highest training performance on Lunar Lander and ranks as the second-best on Car Racing, closely trailing the PPO baseline. Compared to the Initial Policy, which serves as the search starting point, MLES achieves substantial improvements across both tasks, demonstrating its practical effectiveness. These results provide a positive answer to our first research question: MLES is capable of directly synthesizing high-performing policies through environment interaction, delivering performance on par with strong DRL baselines.
@@
Compared to EoH, MLES consistently outperforms it on both training and testing instances. The performance gap is especially notable on the more challenging Car Racing task, indicating that integrating behavior analysis significantly enhances the effectiveness and robustness of evolved policies.
@@
However, we observe a consistent drop in test performance relative to training across all methods. This indicates that generalization remains a challenge, especially in high-variance environments. In Section 4.3, we explore how MLES can leverage its population-based nature to improve generalization through policy ensembling.
Figure 3: Performance convergence on Lunar Lander task
Figure 4: Performance convergence on Car Racing Task
Figure 5: Evolutionary dynamics of Car Racing policies. The plot depicts population score distributions over generations, with yellow lines tracing all ancestors of the best-performing policy. The blue-highlighted lineage is examined in detail to reveal the stepwise improvements guided by IBE-driven insights.
Analysis of policy discovery efficiency
@@
Figs. 3 and 4 illustrate the convergence of the best policy performance with respect to the count of environment resets for four methods. Colored curves represent the mean performance over five runs, and shaded areas indicate the standard error of the mean, reflecting the stability of each method during the policy discovery process. MLES consistently demonstrates superior discovery efficiency compared to both DRL methods (DQN, PPO) and the LES-based baseline (EoH). On the simpler Lunar Lander task, MLES discovers high-performing policies within approximately 5,000 environment resets, which is substantially faster than both PPO and DQN. On the more challenging Car Racing task, MLES achieves a convergence speed comparable to PPO while exhibiting significantly lower variance across runs, indicating more stable learning dynamics. In comparison to EoH, MLES yields both faster convergence and higher final performance across tasks. The narrower confidence intervals in later stages further emphasize MLES’s robustness and consistency. These results highlight the benefits of incorporating visual feedback-driven behavior analysis into the evolutionary process. By leveraging richer signals beyond scalar performance metrics, MLES more effectively guides policy search, leading to faster convergence and enhanced stability.
@@
Another noteworthy point is that both MLES and EoH exhibit higher initial performance compared to DRL baselines. This advantage arises from the ability of the LES paradigm to easily use prior knowledge as initial policies, allowing for a more informed starting point. In contrast, defining meaningful policies in the form of neural networks is challenging, and DRL methods typically rely on randomly initialized weights. This highlights the inherent capacity of MLES to reuse and transfer knowledge. This capability is particularly valuable in domains where expert heuristics are available, enabling more sample-efficient exploration from the outset.
Qualitative analysis of the interpretability

### v2

@@ 2.4 Evolutionary Search for Discovering Policy
This subsection details the evolutionary search loop in MLES. Building upon the EoH framework (Liu et al., 2024a), we integrate behavioral evidence analysis, thereby introducing a prototype implementation of MLES designed to discover high-performing programmatic policies. The pseudocode of MLES is shown in Algorithm 1. At each search step, parent selection determines the starting points, while evolutionary operators guide the generation of new policies. By combining carefully designed operators with principled selection strategies, MLES achieves a dynamic balance between exploration and exploitation, facilitating the search for high-performing policies within the policy space.
Evolutionary operators. The core of MLES is a set of evolutionary operators that drive the generation of new policies. We define four operators, categorized into two types:
@@
  • Multimodal modification operators (M1_M and M2_M): These operators instruct MLLMs to analyze BE and refine existing policies through detailed, informed modification. M1_M asks the MLLMs to identify behavioral shortcomings by jointly analyzing the policy code and its BE, then revise the control logic accordingly. M2_M prompts the MLLMs to identify critical parameters in the policy and adjust them based on the observed evidence, leading to targeted optimization. The integration of BE analysis provides grounded insights for the MLLM’s reasoning, resulting in more coherent and purposeful offspring generation.
Algorithm 1 Pseudocode of MLES
1:Task description, training instances, interfaces of Evaluator and Summarizer
2:Initialize policy pool
3:while termination condition not met do
4:  for each evolutionary operator do
5:   Select parent policies from policy pool
6:   Obtain BEs of parent policies
7:   Construct multimodal few-shot prompt
8:   Generate offspring policy via MLLM
9:   Evaluate offspring and create its BE
10:   Manage and update policy pool
11:  end for
12:end while
13:return A pool of control policies
Parent selection. For each search step, parents are selected from the policy pool using exponential rank selection, a standard method in EC (Blickle & Thiele, 1996). In this scheme, policies are ranked according to their performance, and the selection probability $p_{i}$ for the policy at rank $r_{i}$ is proportional to an exponential decay of its rank, typically expressed as $p_{i}\propto e^{-c\cdot r_{i}}$. This method provides a robust way to control selection pressure, effectively balancing the exploitation of high-performing policies while maintaining diversity by ensuring lower-ranked individuals still have a chance of being selected.
Few-shot prompt construction. After selecting the operator and parents, the prompt sampler constructs a task-specific few-shot prompt to guide the MLLMs in generating new policies. Each prompt consists of the following components: (1) a task description, (2) the code and thought of each parent policy, (3) relevant BEs, and (4) operator-specific instructions. Together, these elements provide the necessary context and guidance for effective policy generation.
Policy generation & individual construction. Guided by the constructed prompts, the LLM ensemble generates a candidate policy. The policy is then executed by evaluators to assess the quantitative metrics. The raw data from its executions is processed by the behavior summarizer, converting it into BE, and ultimately leading to the construction of the corresponding individual.
Policy pool management. Offspring individuals are filtered for redundancy and ranked based on their quantitative metrics. The top $N$ individuals, based on these rankings, are then added back into the policy pool, facilitating progressive refinement and sustained diversity over successive iterations.
@@ 3.1 Experimental Setting
Benchmarks. We evaluate the proposed MLES framework on two representative control tasks from the OpenAI Gym suite (Brockman et al., 2016): Lunar Lander and Car Racing. These tasks collectively cover both discrete and continuous control settings, providing a comprehensive testbed for our method.
  • Lunar Lander is a discrete-action control task where the agent must land a spacecraft on a designated pad. The agent has four action choices, and the environment has an 8-dimensional state space. It is commonly used for benchmarking RL methods in discrete action domains.
@@
Baselines. We compare MLES against both DRL methods and an existing LES approach for direct control policy discovery:
  • Deep Q-Network (DQN) (Mnih et al., 2013) is a value-based DRL algorithm designed for discrete action spaces. It serves as a representative DRL baseline for Lunar Lander.
  • Proximal Policy Optimization (PPO) (Schulman et al., 2017) is a widely used on-policy policy gradient method that can handle both discrete and continuous control tasks.
  • EoH (Liu et al., 2024a) is a recent LES framework for automated code discovery. We use it as a baseline to assess the added value of behavioral evidence analysis in MLES, as it shares the same evolutionary framework but lacks the visual feedback-driven behavior analysis used in MLES.
Beyond the baselines listed above, we also considered other paradigms for generating human-readable policies, such as genetic programming (GP) and policy distillation. However, GP-based approaches are highly dependent on manually-designed DSLs (Trivedi et al., 2021; Qiu & Zhu, 2022). Adapting them to our benchmarks would require non-trivial effort, making a fair comparison difficult. On the other hand, distillation-based methods typically incur a performance loss compared to their DRL teacher models (Verma et al., 2018). Therefore, evaluating MLES against high-performing DRL methods is the most direct and meaningful measure of its capabilities.
Implementation details. Both MLES and EoH employ GPT-4o-mini for policy generation. The LLM query budget is set to 2000 requests, corresponding to 10,000 and 8,000 environment resets for Lunar Lander and Car Racing, respectively. To ensure a fair comparison, we train all baselines under an identical budget of environment resets. The policy pool size $N$ is set to 16, with the initial population of policies for MLES shared with EoH to avoid any biases arising from different starting points. All experiments are conducted on a machine equipped with an Intel Core i9-13980HX processor and 32 GB of RAM. For the Lunar Lander task, we select five representative instances for training, while for the Car Racing task, four representative tracks are chosen. Both tasks use ten test environments, each generated from a random seed in the range zero to nine. Each experiment is repeated five times to ensure the robustness and reliability of the results.
@@ 3.2 Experimental Results
Table 1: Performance on benchmarks. Best results are in bold, second-best are underlined.
@@
Training Testing Training Testing
Average SEM Best run Average Average SEM Best run Average
DQN 1.017 $\pm$0.022 1.066 0.508 79.777 $\pm$2.635 91.182 71.724
PPO 1.032 $\pm$0.026 1.076 0.846 99.212 $\pm$0.390 100.000 94.546
@@
Eoh 1.053 $\pm$0.021 1.085 0.776 89.808 $\pm$1.553 96.675 79.286
MLES 1.090 $\pm$0.005 1.098 0.819 98.070 $\pm$0.719 100.000 96.358
Comparative evaluations of MLES. Table 1 presents the performance of each method on Lunar Lander and Car Racing benchmarks. We report the average and best performance from five independent runs for both training and testing instances, along with the standard error of the mean (SEM). On the Lunar Lander task, a score above 1.00 signifies a successful landing on all instances, with higher scores indicating greater fuel savings. For the Car Racing task, a score of 100 represents a flawless completion of all tracks. For both tasks, a higher score indicates better performance. The specific metrics used for these evaluations are detailed in Appendix C.
@@
As Table 1 shows, the MLES achieves the highest training performance on Lunar Lander and ranks second-best on Car Racing, performing comparably to the strong PPO baseline. Compared to the Initial Policy, which serves as our starting point, MLES demonstrates substantial improvements across both tasks, highlighting its practical effectiveness. These results provide strong evidence that MLES is capable of synthesizing high-performing policies directly through environment interaction, delivering performance on par with strong DRL baselines.
@@
When compared to EoH, MLES consistently and significantly outperforms it on both training and testing instances. Furthermore, MLES exhibits superior convergence and stability, as indicated by its smaller SEM. The performance gap is especially pronounced in the more challenging Car Racing task. This outcome demonstrates that integrating behavior analysis not only substantially enhances the effectiveness of policy evolution but also improves the robustness of the generated policies.
@@
However, we observe a consistent drop in testing performance relative to training across all methods. This indicates that generalization remains a challenge. In Appendix B.5, we explore how MLES can leverage its population-based nature to improve generalization through policy ensembling.
Analysis of policy discovery efficiency. Figs. 4 and 4 illustrate the convergence of the best policy performance with respect to the count of environment resets for four methods. Colored curves represent the mean performance over five runs, and shaded areas indicate the SEM, reflecting the stability during the policy discovery process.
Figure 3: Convergence on Lunar Lander task
Figure 4: Convergence on Car Racing task
@@
As demonstrated, the MLES consistently shows superior search efficiency compared to the baselines. On the simpler Lunar Lander task, MLES discovers high-performing policies within approximately 5,000 environment resets, which is substantially faster than both PPO and DQN. On the more challenging Car Racing task, MLES achieves a convergence speed comparable to PPO while exhibiting significantly lower variance across runs, indicating more stable learning dynamics. In comparison to EoH, MLES achieves both faster convergence and higher final performance across tasks. The narrower confidence intervals in later stages further emphasize MLES’s robustness and consistency. These results highlight the benefits of incorporating visual feedback-driven behavior analysis into the evolutionary process. By leveraging richer signals beyond scalar performance metrics, MLES more effectively guides policy search, leading to faster convergence and enhanced stability.
@@
Another noteworthy observation is that both MLES and EoH exhibit higher initial performance compared to DRL baselines. This advantage stems from the ability of the LES paradigm to easily use prior knowledge as initial policies, allowing for a more informed starting point for the search. In contrast, defining meaningful policies in the form of neural networks is challenging, and DRL methods typically rely on randomly initialized weights. This highlights the inherent capacity of MLES to reuse and transfer knowledge. This capability is particularly valuable in domains where expert heuristics are available, enabling more sample-efficient exploration from the outset.